(Carla) Bell Industries sells samsung phones for CDN$100. The unit variable cost per phone is $50...

(Carla)

Bell Industries sells samsung phones for CDN$100.
The unit variable cost per phone is $50 plus a selling commission of 10%. Fixed manufacturing costs total $1,250
per month, while fixed selling and administrative costs total $2,500.
a. What is the contribution margin per phone?
Seling Price - Variable Cost
CND$100 - $50 = $50.00
b. What is the breakeven point in phones?
Fixed cost ÷ Contribution Margin per phone
$1,250 + $2,500 = $3,750.00 ÷ $50.00 = 75 phones
c. How many phones must be sold to earn a targeted profit of $7,500
Target Profit = Fixed Cost + Target Profit ÷ Contribution Margin

$3,750 + $7,500 = $11,250 ÷ 50 = 225

Are these answers correct/wrong. Comment on answers.
Answer #1

A.Contribution margin per phone = Selling price - variable cost - Commission

=100-50 - (10%×50)= 100-50-10

= 40

B. Break even points = fixed cost/ contribution margin

=(1250+2500)/40= 3750/40

=93.75 or 94

C. No of phone to be sold earn Target profit = Fixed cost + Target profit /contribution margin

= (3750+7500)/40= 11250/40

= 281.25 or 282